OpenStudy (aaliyahhailey):

a 25ft ladder is leaning against a building. the base of the ladder is 6ft away from the building. how high up is the ladder?

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

do you agree with the drawing?

OpenStudy (aaliyahhailey):

yes!

jimthompson5910 (jim_thompson5910):

have you learned the pythagorean theorem?

OpenStudy (aaliyahhailey):

i have! a(2) + b(2) = c(2) correct?

jimthompson5910 (jim_thompson5910):

I think you meant to say this? \[\Large a^2 + b^2 = c^2\]

jimthompson5910 (jim_thompson5910):

typing that on the keyboard, you'd say `a^2+b^2=c^2`

OpenStudy (aaliyahhailey):

yeah that's what i meant LOL (it's harder for me on a mac cause idk how to do the 2nd power option)

jimthompson5910 (jim_thompson5910):

you hold the shift key down, then you hit the "6" key http://prntscr.com/b9nyu4 to get the `^` symbol

OpenStudy (aaliyahhailey):

ah alright! okay, i just need help starting off the problem and i'll understand the rest!

jimthompson5910 (jim_thompson5910):

anyways, using the pythagorean theorem gives us \[\Large a^2 + b^2 = c^2\] \[\Large x^2 + 6^2 = 25^2\] \[\Large x^2 + 36 = 625\] how do we solve for x?

OpenStudy (aaliyahhailey):

wouldn't you use the 25^2 + 36?

jimthompson5910 (jim_thompson5910):

in this case, a = x b = 6 where 'a' and 'b' are the legs c = 25 is the hypotenuse

jimthompson5910 (jim_thompson5910):

IF 25 and 6 were the two legs, then you'd have 25^2+6^2 = 625+36 but that's not the case here

jimthompson5910 (jim_thompson5910):

make sense? or no?